OpenStudy (anonymous):

scott is 4 times older than beth. In six years from now scott will be 3 times older than beth. Write down 2 equations to discribe the relationship between there ages and solve the equations and how old are both of them????

OpenStudy (anonymous):

if beth is x years old now, how old is scott now?

OpenStudy (anonymous):

how old are both of them

OpenStudy (anonymous):

I am asking you. suppose beth is B years old. how old is scott now?

OpenStudy (anonymous):

it does not say

OpenStudy (anonymous):

it does say. read the problem.

OpenStudy (anonymous):

i typed the question how it is in my book

OpenStudy (anonymous):

okay suppose you are beth and I am scott. if your age is 15, what is my age?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

why 60?

OpenStudy (anonymous):

15*4 timesher ages

OpenStudy (anonymous):

okay good. now, suppose your age is B years. how old am I?

OpenStudy (anonymous):

B*4=b^4

OpenStudy (anonymous):

okay. I am four times your age. so if your age is B, then my age is 4 times B. so suppose my age is S years. I can say that my age = 4 times your age or S = 4 x B right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

in six years, what would your age and my age be?

OpenStudy (anonymous):

suppose you are 15 years old now. what will your age be in 6 six years?

OpenStudy (anonymous):

90

OpenStudy (anonymous):

what? so if your age is 15 in 2011 you will be 90 years old in 2017?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how old are you now?

OpenStudy (anonymous):

sorry my mind is running low i will be 21

OpenStudy (anonymous):

okay good. so if you are B years old now, how old would you be 6 years from now?

OpenStudy (anonymous):

6*b

OpenStudy (anonymous):

why 6 times B?

OpenStudy (anonymous):

if you are B years old now, you will be B+6 years old in 6 years.

OpenStudy (anonymous):

sorry yes i understand that part

OpenStudy (anonymous):

similarly I will be S+6 years old in 6 years. by that time my age will only be three times your age that means s+6 = 3 times (b+6)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so s =4b ----------------------> equation 1 s+6 = 3(b+6) ----------------> equation 2 solve the two equations and find s and b

OpenStudy (anonymous):

i cant can you do it and show me please

OpenStudy (anonymous):

http://www.purplemath.com/modules/systlin5.htm

OpenStudy (anonymous):

that is the extent of my help. all the best!